Demo file for DVIWIN 2.0 January 15, 1993 Hippocrates Sendoukas

This driver lets you preview and print any DVI file under MS-Windows 3.x. Its main advantages are: speed, compatibility with most screens and printers, and graphics capability. Apart from some assembly acceleration for all machines, it uses special routines for 80386 or 80486 processors to speed up the rasterization. All screen and printer handling is done through Windows, so it should work on any printer supported by Windows. The graphics capabilities were needed to incorporate graphs in my thesis and I was tired of manual cutting and pasting. In its present state, you can put arbitrary graphics files produced by most Windows application, or other graphics files (eg. TIFF, PCX, etc.) provided that you have the appropriate graphics filter. The following graph was created by Excel:

Sample Figure: residuals vs. maturity

For more information on using the programs, please read or print the files “dviwin.wri”, “wbr.wri” and “clipmeta.wri”. If you encounter any installation or operation problems, please let me know. My e–mail address is “sendouk@scf.usc.edu”.